# Antichain cutsets in real-ranked lattices

### v1

@@ 2. Definitions and background
Von Neumann defines a continuous projective geometry over a field $\mathbb{F}$ [19, Section 5]; see also [1, Chapter X.5]. In the language above, this may be described as follows. Let $L_{n}$ be the lattice of all subspaces of the vector space $\mathbb{F}^{n}$. When $n$ is a multiple of $k$, view $\mathbb{F}^{n}$ as $\mathbb{F}^{k}\oplus\mathbb{F}^{k}\oplus\cdots\oplus\mathbb{F}^{k}$. Now define $\varphi$ by mapping a subspace $W$ of $\mathbb{F}^{k}$ to $W\oplus\cdots\oplus W$. It is easy to see that $\varphi$ preserves renormalized rank, and is a lattice mapping. Every element of the metric completion is rank modular with respect to $\rho$.
@@
The measurable Boolean lattice on $[0,1]$ may be built from this machinery, as follows [20, Section 8]. Let $L_{n}$ be the Boolean lattice of all subsets of $[n]$. When $n$ is a multiple of $k$, view $[n]$ as $k]\times[n/k]$. Define $\varphi$ by mapping $S\subseteq[k]$ to $S\times[n/k]$, and take the directed limit and metric completion.
@@
Haiman gives explicit descriptions of Björner’s measurable partition lattice as well as of related lattices [12]. Haiman’s lattices are again rank supersolvable [12, Lemma 2.15]. A recent preprint of Contreras Mantilla and Sinclair [5] discusses the Von Neumann and Björner construction from a model theory perspective. See also the paper [18] of Vershik and Yakubovich, and the survey article of Björner [3].

### v2

# arXiv:2512.14826v2[math.CO] 01 Mar 2026

@@ 1. Introduction
In contrast, the (full) Boolean lattice on a infinite set admits only the trivial antichain cutsets $\{\hat{0}\}$ and $\{\hat{1}\}$. For a countably infinite set, this is straightforward, as follows. The subposet of all finite subsets forms a (non-complete) sublattice. By countability, adjoining $\hat{1}$ to a maximal chain in this sublattice yields a chain that is maximal in the full Boolean lattice. The subposet of all subsets with a finite complement is dual, and entirely similar. It now follows by the main result of [8] that any putative nontrivial antichain cutset must contain both all sets of some size $n$, and likewise also all sets whose complement has some size $n^{\prime}$. But this clearly contradicts the antichain property! In a recent preprint [11], Ginsburg and Sands have shown the stronger result that any cutset in the Boolean lattice over a set of infinite cardinality $\kappa$ must contain a chain having strictly larger cardinality than $\kappa$.
@@ 2. Definitions and background
Von Neumann defines a continuous projective geometry over a field $\mathbb{F}$ [20, Section 5]; see also [1, Chapter X.5]. In the language above, this may be described as follows. Let $L_{n}$ be the lattice of all subspaces of the vector space $\mathbb{F}^{n}$. When $n$ is a multiple of $k$, view $\mathbb{F}^{n}$ as $\mathbb{F}^{k}\oplus\mathbb{F}^{k}\oplus\cdots\oplus\mathbb{F}^{k}$. Now define $\varphi$ by mapping a subspace $W$ of $\mathbb{F}^{k}$ to $W\oplus\cdots\oplus W$. It is easy to see that $\varphi$ preserves renormalized rank, and is a lattice mapping. Every element of the metric completion is rank modular with respect to $\rho$.
@@
The measurable Boolean lattice on $[0,1]$ may be built from this machinery, as follows [21, Section 8]. Let $L_{n}$ be the Boolean lattice of all subsets of $[n]$. When $n$ is a multiple of $k$, view $[n]$ as $k]\times[n/k]$. Define $\varphi$ by mapping $S\subseteq[k]$ to $S\times[n/k]$, and take the directed limit and metric completion.
@@
Haiman gives explicit descriptions of Björner’s measurable partition lattice as well as of related lattices [13]. Haiman’s lattices are again rank supersolvable [13, Lemma 2.15]. A recent preprint of Contreras Mantilla and Sinclair [5] discusses the Von Neumann and Björner construction from a model theory perspective. See also the paper [19] of Vershik and Yakubovich, and the survey article of Björner [3].